OpenStudy (anonymous):

What happens when you add two square roots? For example root of 3 plus root of 3

OpenStudy (anonymous):

\[\sqrt{3}+\sqrt{3}\]

OpenStudy (earthcitizen):

you get \[2\sqrt{3}\]

hartnn (hartnn):

\(\sqrt a+\sqrt a=2\sqrt a\)